# thanks: Haihao Shi and Zhenyang Huang contribute equally to this workthanks: Corresponding authorthanks: Corresponding author

### v1

@@ IV Sparse Identi Cation of Nonlinear system part
Although KANs can successfully build analytical network for the neutrino fast flavor conversion, we identified two key limitations: the resulting formulas can lack the desired simplicity, and the process of symbolic regression sometimes struggles with randomness-induced bias. To overcome these specific challenges and to construct a more concise, low-rank model, we employ the Sparse Identification of Nonlinear Dynamics (SINDy) framework to extract interpretable models from input data. SINDy assumes that the system dynamics can be described by a sparse combination of candidate nonlinear functions. The system is modeled as a general nonlinear dynamical system [72]:
@@
$\hat{\bm{\xi}}_{k}=\arg\min_{\bm{\xi}^{\prime}_{k}}||\dot{\mathbf{X}}_{k}-\bm{\Theta}(\mathbf{X})\bm{\xi}^{\prime}_{k}||_{2}+\alpha||\bm{\xi}^{\prime}_{k}||_{2}\text{,}\\ \text{for each column }k\text{ of }\dot{\mathbf{X}}\text{,}$ (26)
@@
$\mathcal{S}_{k,i}=\sqrt{(\mu^{*}_{k,i})^{2}+(\sigma_{k,i})^{2}}\text{}$
@@
$\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{8}\end{array}\right]=\mathbf{A}\left[\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{11}\end{array}\right]+\mathbf{B}\left[\begin{array}[]{c}\sin\left(x_{1}\right)\\ \sin\left(x_{2}\right)\\ \vdots\\ \sin\left(x_{11}\right)\end{array}\right]+\mathbf{C}$ (27)
@@ V Neutrino Energy Spectra Reconstruction and Comparative Analysis
$L_{\nu_{e}}:L_{\bar{\nu}_{e}}:L_{\nu_{x}}$ $=1:1:0.33,$ (30)
$\langle E_{\nu_{e}}\rangle:\langle E_{\bar{\nu}_{e}}\rangle:\langle E_{\nu_{x}}\rangle$ $=9:12:16.5,$
$\eta_{\nu_{e}}:\eta_{\bar{\nu}_{e}}:\eta_{\nu_{x}}$ $=3.2:4.5:2.3.$
@@
The results illustrate that no single model achieves uniformly superior predictive accuracy across all neutrino species. The Parametric Kolmogorov-Arnold Network (Param. KAN) demonstrates the lowest prediction errors for the $\nu_{e}$ and $\nu_{x}$ species, while the baseline MLP model achieves the smallest errors for $\bar{\nu}_{e}$ and $\bar{\nu}_{x}$. The Symbolic KAN yields moderate errors overall, performing better than the PySINDy-based methods. In contrast, the PySINDy and PySINDy-SA models, despite their advantages in interpretability—with PySINDy-SA being the most interpretable—consistently show the highest prediction errors in this comparison.
@@
From the perspective of model architecture and interpretability, these methods span a spectrum from “black-box” to “white-box.” The MLP is a classic black-box model, with opaque internal logic. At the other extreme, the PySINDy methods serve as white-box models, distilling system dynamics into concise ordinary differential equations (ODEs) that are ideal for uncovering physical laws. Notably, PySINDy-SA is specifically designed to deliver the most interpretable, concise expressions. The KAN model occupies an intermediate position: its parametric form achieves high-fidelity predictions, while its symbolic form can generate interpretable formulas. It is important to note, however, that these symbolic formulas are often highly complex.
@@
Taken together, our analysis highlights a fundamental trade-off in data-driven scientific modeling: balancing predictive accuracy with physical interpretability. If the primary goal is to achieve the highest possible accuracy in predicting system states, models like the Param. KAN or MLP are advantageous. Conversely, for research focused on discovering and understanding fundamental physical laws, methods such as PySINDy and PySINDy-SA—which produce concise, interpretable governing equations—are indispensable, even if this comes at the cost of reduced predictive accuracy. While Symbolic KAN offers a potential pathway for unifying high accuracy with interpretability, its resulting complexity suggests that truly interpretable models must also remain sufficiently concise to yield human-understandable, meaningful physical insight.
@@
Table 2: Comparison of different models for reconstructing post-FFC supernova neutrino energy spectra. The models are evaluated on their reconstruction accuracy, quantified by the spectral shape error $\delta_{\nu}$ for each neutrino ($\nu_{e}$, $\bar{\nu}_{e}$, $\nu_{x}$, $\bar{\nu}_{x}$). For each method, the values presented correspond to the best-performing result obtained from multiple training runs or configurations. Lower $\delta_{\nu}$ indicates higher accuracy. We also report an overall average accuracy, defined as $(1-\overline{\delta_{\nu}})\times 100\%$, where $\overline{\delta_{\nu}}$ is the mean of the $\delta_{\nu}$ values across the four neutrino types. Additionally, model form and interpretability are also described.
Model $\delta_{\nu_{e}}$ $\delta_{\bar{\nu}_{e}}$ $\delta_{\nu_{x}}$ $\delta_{\bar{\nu}_{x}}$ Average Accuracy Model Form & Interpretability
MLP [67] 0.20 0.06 0.23 0.05 86% Black-box neural network (High Accuracy)
Param. KAN 0.11 0.08 0.13 0.10 90% Parametric B-splines (High Accuracy)
Symbolic KAN 0.20 0.18 0.13 0.16 83% Complex symbolic formula (Interpretable)
PySINDy 0.21 0.21 0.19 0.19 80% Concise ODEs (Parsimonious and Interpretable)
PySINDy-SA 0.23 0.21 0.19 0.19 80% Concise ODEs (Most Parsimonious and Interpretable)
@@
Figure 4: This figure compares the reconstructed post-FFC energy spectra from our models with exact calculations. The exact calculations are performed using the benchmark setup described in [67]. The panels show results for $\nu_{e}$ (top-left), $\bar{\nu}_{e}$ (top-right), $\nu_{x}$ (bottom-left), and $\bar{\nu}_{x}$ (bottom-right). The legend indicates results from Exact (blue squares), Parametric KAN (red circles), Symbolic KAN (orange circles), and PySINDy (green circles).
@@ VI Data-driven Exploration of Physical Information
As shown in Figure 2, despite employing diverse machine learning training methodologies, our converged results consistently indicate that the parameter $F_{\text{total}}$ is not utilized in predicting the final neutrino momentum. However, satisfactory outcomes cannot be achieved if it’s removed from the dataset prior to training. This suggests that $F_{\text{total}}$ is still involved in the training process; however, its spectral dependence likely approximates a constant relationship rather than a variable one, leading to its simplification by the network in the final stages. This aligns with the previously proposed theory that in a multi-energy neutrino gas, when the FFC growth rate significantly surpasses that of the vacuum Hamiltonian, all neutrinos exhibit a common survival probability determined by the energy-integrated spectrum (regardless of energy) [67]. However, Figure 2 shows that the energy input still has a small weight, which is possibly due to energy-dependent neutrino emission and absorption occurring later in the evolution [78]. Utilizing the methodology described above, this study obtains the analytical and semi-analytical dependence of the energy spectrum on its moments. Formally, the dependencies derived from both KAN and SINDy can be represented by matrix equations of an identical form. This may indicate that for most learning tasks, there exists an underlying, well-performing, low-rank neural network solution [79].
While KAN can yield symbolic expressions, their complexity makes them less suitable for extracting physical insights than the more concise SINDy-based methods. Table 2 show that the SINDy and SINDy-SA models have nearly identical predictive accuracy. However, since SINDy-SA is specifically designed to produce the most simplified form, it represents the ideal candidate for our analysis. Consequently, this section will focus exclusively on the insights derived from the SINDy-SA model.
To identify the key input physical variables influencing the FFC outputs, we computed the Mutual Information (MI) based on the results obtained from SINDy-SA. MI is an information-theoretic metric that quantifies the amount of information shared between two random variables. Formally, the MI between variables $X$ and $Y$ is defined as [80]:
$I(X;Y)=\iint p(x,y)\log\frac{p(x,y)}{p(x)p(y)}dxdy.$
Importantly, it can detect both linear and nonlinear dependencies without assuming any specific functional form. In the context of feature selection or sensitivity analysis, MI provides a model-agnostic way to evaluate the relevance of input variables for predicting outputs. The MI heatmap is shown in Figure 5. To complement the mutual information analysis, we also employed SHapley Additive exPlanations (SHAP) [81, 82] to investigate the input-output relationships from a model-based perspective. The feature importance rankings derived from SHAP proved to be robust across different explainers (LightGBM, RandomForest, and XGBoost) and strongly corroborated the findings presented in Figure 5. Given this consistency, a detailed discussion is omitted for brevity. The complete code is available at the link provided in Code and Data Availability.
Across all input parameters, $n_{\nu_{x}}$ is observed to be the most influential and exhibits the strongest correlation with all outputs. This is not a surprising result, as it has a similar structure to previous work showing that the frequency of collective oscillations scales as $\sim\sqrt{\omega\mu_{\nu}}$ [83], where $\omega$ is the vacuum neutrino oscillation frequency and $\mu_{\nu}=\sqrt{2}G_{F}n_{\nu}$ represents the potential due to a neutrino density $n_{\nu}$. In particular, this indicates that in the specific astrophysical model generating our dataset, reactions producing heavy-lepton neutrinos (e.g., electron–positron pair annihilation) influence the results of Fast Flavor Conversions more significantly than those involving electron-type neutrinos.
One potential explanation for our finding-that $n_{\nu_{x}}$ is the most influential parameter -may be related to the unique angular distribution of $\nu_{x}$. The fact that its angular spread is the smallest among all flavors [84] could provide favorable conditions for the $\nu$ELN crossing required to trigger FFCs, thus offering a possible physical basis for our model’s result. Building on this, one could further speculate on a potential feedback loop: once triggered, FFCs could increase the luminosity of these same heavy-lepton neutrinos. This enhancement could then be one of the factors contributing to nontrivial modifications in the supernova’s energy transport [32, 85].
Figure 5: Heatmap of the MI scores quantifying the statistical dependence between input physical variables and output moments. Ml is a robust measure that captures both linear and non-linear relationships. The scores are plotted on a logarithmic color scale, where warmer colors (red) indicate stronger dependence and cooler colors (blue) signify weaker dependence. The analysis prominently identifies the $n_{\nu_{x}}$ as the most influential input variable, exhibiting the highest degree of correlation with the output moments, particularly $I_{0}$.
@@
From the MI heatmap computed using the SINDy-SA framework, we find that among the many input parameters determining the asymptotic values of the neutrino radial moments $I_{0}$ and $I_{1}$, only the initial number density of heavy-lepton neutrinos exhibits significant correlation.The mutual information values for all other parameters are negligible. It is important to note that most of these parameters in fact are the initial-state combinations of the neutrino radial moments prior to the onset of FFCs. An interpretation was previously provided at the beginning of this section using the KAN model analysis, and here the mutual information approach—capable of capturing both linear and nonlinear dependencies—arrives at the same conclusion. This consistency may prompt us to consider the problem from a new perspective: while the triggering of FFCs is known to be dependent on the details of the initial angular distribution, the connection between the asymptotic outcome and those specific initial conditions weakens rapidly once FFCs are triggered, and the system’s dynamics become dominated by macroscopic properties such as the global particle number density.
@@
This behavior shares a conceptual similarity with the well-known “sandpile model” in dynamical systems exhibiting self-organized criticality [86]: the triggering of an avalanche may be caused by a single, randomly falling grain of sand, but the size, shape, and duration of the avalanche are almost entirely independent of the properties of that grain, instead being determined by the sandpile’s overall “critical state” , such as its slope and density. Once the sandpile reaches this critical state, there is no longer any correlation between the system’s response to a perturbation and the details of that perturbation itself. Similarly, in our scenario, specific initial combinations of $I_{0}$ and $I_{1}$ act as the triggering grain of sand, serving mainly to determine whether the FFC “avalanche” is initiated. Once triggered, the system’s nonlinear dynamics quickly “forget” the details of the trigger, and the final asymptotic state is governed by the system’s global macroscopic properties—in this study, represented by the heavy-lepton neutrino number density. It should be noted that this agrees with the results of [67], while offering a view of the underlying dynamics.

### v2

@@ IV Sparse Identi Cation of Nonlinear system part
Although KANs can successfully build analytical network for the neutrino fast flavor conversion, we identified two key limitations: the resulting formulas remain relatively complex, and the process of symbolic regression sometimes struggles with randomness-induced bias. To overcome these specific challenges and to construct a more concise, low-rank model, we employ the Sparse Identification of Nonlinear Dynamics (SINDy) framework to extract interpretable models from input data. SINDy assumes that the system dynamics can be described by a sparse combination of candidate nonlinear functions. The system is modeled as a general nonlinear dynamical system [76]:
@@
[…]
@@
$\mathcal{S}_{k,i}=\sqrt{(\mu^{*}_{k,i})^{2}+(\sigma_{k,i})^{2}}\text{}.$
@@
$\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{8}\end{array}\right]=\mathbf{A}\left[\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{11}\end{array}\right]+\mathbf{B}\left[\begin{array}[]{c}\sin\left(x_{1}\right)\\ \sin\left(x_{2}\right)\\ \vdots\\ \sin\left(x_{11}\right)\end{array}\right]+\mathbf{C}.$ (27)
@@ V Neutrino Energy Spectra Reconstruction and Comparative Analysis
[…]
@@
The results illustrate that no single model achieves uniformly superior predictive accuracy across all neutrino species. The Parametric Kolmogorov-Arnold Network (Param. KAN) demonstrates the lowest prediction errors for the $\nu_{e}$ and $\nu_{x}$ species, while the baseline MLP model achieves the smallest errors for $\bar{\nu}_{e}$ and $\bar{\nu}_{x}$. The Symbolic KAN yields moderate errors overall, performing better than the SINDy-based methods. In contrast, the SINDy and SINDy-SA models, despite their advantages in interpretability—with SINDy-SA being the most interpretable—consistently show the highest prediction errors in this comparison.
@@
From the perspective of model architecture and interpretability, these methods span a spectrum from “black-box” to “white-box.” The MLP is a classic black-box model, with opaque internal logic. At the other extreme, the SINDy methods serve as white-box models, distilling system dynamics into concise ordinary differential equations (ODEs) that are ideal for uncovering physical laws. Notably, SINDy-SA is specifically designed to deliver the most interpretable, concise expressions. The KAN model occupies an intermediate position: its parametric form achieves high-fidelity predictions, while its symbolic form can generate interpretable formulas. It is important to note, however, that these symbolic formulas are often highly complex.
@@
Taken together, our analysis highlights a fundamental trade-off in data-driven scientific modeling: balancing predictive accuracy with physical interpretability. Our results exemplify this: if the primary goal is to achieve the highest possible accuracy in predicting system states, models like the Param. KAN or MLP are advantageous. Conversely, for research focused on discovering and understanding fundamental physical laws, methods such as SINDy and SINDy-SA—which produce concise, interpretable governing equations—are indispensable, even if this comes at the cost of reduced predictive accuracy. What scientists need to do is not a simple compromise, but a strategic choice from a collection of high-performing models. While Symbolic KAN offers a potential pathway for unifying high accuracy with interpretability, its resulting complexity suggests that truly interpretable models must also remain sufficiently concise to yield human-understandable, meaningful physical insight.
@@
Table 2: Comparison of different models for reconstructing post-FFC supernova neutrino energy spectra. The models are evaluated on their reconstruction accuracy, quantified by the spectral shape error $\delta_{\nu}$ for each neutrino ($\nu_{e}$, $\bar{\nu}_{e}$, $\nu_{x}$, $\bar{\nu}_{x}$). For each method, the values presented correspond to the best-performing result obtained from multiple training runs or configurations. Lower $\delta_{\nu}$ indicates higher accuracy. We also report an overall average accuracy, defined as $(1-\overline{\delta_{\nu}})\times 100\%$, where $\overline{\delta_{\nu}}$ is the mean of the $\delta_{\nu}$ values across the four neutrino types. Additionally, model form and interpretability are also described. We use SymPy’s count_ops function to compute the complexity metric; by counting operations across nested subexpressions, it captures both operation quantity and effective depth, providing a structural complexity indicator. Parame. KAN denotes the best-performing test result of the original KAN with learned B-spline activations (no symbolic regression). Symbolic KAN denotes the symbolic results obtained by applying the symbolic regression procedure (mentioned at the end of section III) to the trained KAN.
Model $\delta_{\nu_{e}}$ $\delta_{\bar{\nu}_{e}}$ $\delta_{\nu_{x}}$ $\delta_{\bar{\nu}_{x}}$ Average Accuracy Model Form & Complexity
MLP [65] 0.20 0.06 0.23 0.05 86% Black-box neural network (Too many)
Param. KAN 0.11 0.08 0.13 0.10 90% Parametric B-splines (Too many)
Symbolic KAN 0.20 0.18 0.13 0.16 83% Complex symbolic formula (1624)
SINDy 0.21 0.21 0.19 0.19 80% Concise ODEs (456)
SINDy-SA 0.23 0.21 0.19 0.19 80% Concise ODEs (295)
@@
Figure 4: This figure compares the reconstructed post-FFC energy spectra from our models with exact calculations. The exact calculations are performed using the benchmark setup described in [65]. The panels show results for $\nu_{e}$ (top-left), $\bar{\nu}_{e}$ (top-right), $\nu_{x}$ (bottom-left), and $\bar{\nu}_{x}$ (bottom-right). The legend indicates results from Exact (blue squares), Param. KAN (red circles), Symbolic KAN (orange circles), and SINDy (green circles).
@@ VI Data-driven Exploration of Physical Information
Compared with conventional data-driven neural networks, KAN and SINDy provide symbolic models rather than black-box predictors. This enables direct examination of the relationship between inputs and outputs, that is, how each input variable affects the predicted results. A practical way to study this relationship is to compute partial derivatives. This allows us to see, in an intuitive manner, how every output responds to each input. The same idea has been used in symbolic KAN-based convolutional neural networks [74]. In our analysis, we compute the partial derivative of each output with respect to each input, then evaluate these derivatives at the original data points, and take absolute values. Table 2 shows that the SINDy and SINDy-SA models have almost identical predictive accuracy. However, because SINDy-SA is specifically designed to yield the most simplified form, it is the ideal candidate for our analysis. Therefore, this section will focus solely on the insights gained from the KAN and SINDy-SA models. The resulting heat map is shown in Figure 5 and Figure 6. We interpret the normalized magnitude as the contribution of each variable to the outcome, where larger values indicate stronger contributions.
Figure 5: Heatmap of normalized absolute partial derivatives for the SINDy-SA. Each entry reports the sample mean of $|\partial y_{j}/\partial x_{i}|$ evaluated at the original data points Warmer colors indicate stronger influence of input variable $x_{i}$ on output $y_{j}$, cooler colors indicate weaker influence. Here, we set all values less than $10^{-3}$ to zero.
Figure 6: Heatmap of normalized absolute partial derivatives for the symbolic KAN. Each entry reports the sample mean of $|\partial y_{j}/\partial x_{i}|$ evaluated at the original data points Warmer colors indicate stronger influence of input variable $x_{i}$ on output $y_{j}$, cooler colors indicate weaker influence. Here, we set all values less than $10^{-3}$ to zero.
@@
As shown in Figure 2 and Figure 6, despite employing diverse machine learning training methodologies, our converged results consistently indicate that the parameter $F_{\text{total}}$ is not utilized in predicting the final neutrino momentum. However, satisfactory outcomes cannot be achieved if it’s removed from the dataset prior to training. This suggests that $F_{\text{total}}$ is still involved in the training process; a possible explanation is that its spectral dependence likely approximates a constant relationship rather than a variable one, leading to its simplification by the network in the final stages. This aligns with the previously proposed theory that in a multi-energy neutrino gas, when the FFC growth rate significantly surpasses that of the vacuum Hamiltonian, all neutrinos exhibit a common survival probability determined by the energy-integrated spectrum (regardless of energy) [65]. However, Figure 2 shows that the energy input ($E_{\nu_{e},i}$ and $E_{\bar{\nu}_{e},i}$) still has a weak influence on the predictions of the output. This is consistent with the statement in [65] that ”the survival probabilities are nearly independent of the neutrino energy.” While this does not confirm complete energy independence, it implies that any existing dependence is weak and non-dominant, potentially stemming from subtle, underlying physical processes. One explanation is that the growth rate of FFC is too rapid compared to the typical timescale of the vacuum term, making the explicit energy dependence in the vacuum term negligible. In this sense, FFC is almost independent of energy, though not entirely. In particular, when energy-independent flavor conversion from fast flavor instability reaches nonlinear saturation, the energy contribution from the vacuum term could appear, including slow flavor conversion. For SINDy-SA, since it is inherently an effective model that seeks low rank, its ability to capture details is lower than that of KAN, which is why its prediction accuracy is lower than that of KAN. However, it can still effectively capture the dominant components in the mapping relationship. From Figure 5, it can also be observed that the dominant parameters for the predicted output are $n_{\nu_{x}}$ and $F_{\nu_{x},i}$, which is consistent with the conclusions we obtained from KAN. Utilizing the methodology described above, this study obtains the analytical and semi-analytical dependence of the energy spectrum on its moments. Formally, the dependencies derived from both KAN and SINDy can be represented by matrix equations of an identical form. This may indicate that for most learning tasks, there exists an underlying, well-performing, low-rank neural network solution [82].
Across all input parameters, the input related to $\nu_{x}$ ($n_{\nu_{x}}$ and $F_{\nu_{x},i}$) is observed to be the most significant and exhibits the strongest correlation with all outputs. Especially, $n_{\nu_{x}}$ consistently holds a prominent position across different models; this is not a surprising result. Because it has a similar structure to previous work showing that the frequency of collective oscillations scales as $\sim\mu_{\nu}$ [83], where $\mu_{\nu}=\sqrt{2}G_{F}n_{\nu}$ represents the potential due to a neutrino density $n_{\nu}$. Furthermore, because $F_{\nu,i}$ is defined as $F_{\nu,i}=(I_{1}/I_{0})_{\nu,i}$ with $I_{0,i}=n_{\nu,i}$, $F_{\nu_{x},i}$ has an explicit dependence on $n_{\nu_{x}}$. This dependence could help account for the relatively large weight of $F_{\nu_{x},i}$ in the heat map. Still, this should be read as a correlation rather than proof of causal dominance, since $F_{\nu_{x},i}$ also reflects $I_{1,\nu_{x},i}$ and model-specific interactions among inputs. Our analysis indicates that reactions affecting the initial number density of heavy-lepton neutrinos, such as electron-positron pair annihilation, play a more pivotal role in shaping the outcome of FFCs within our model.This is because the system’s evolution exhibits the highest sensitivity to the initial number density of heavy-lepton neutrinos, far exceeding its sensitivity to the electron neutrino density or other physical parameters. From another perspective, through flavor conversion, the energy spectra of heavy-leptonic flavors are converted to the electron-type. In this sense, it seems normal that $\nu_{x}$ makes a significant contribution. By contrast, it is strange that the heatmap of $\nu_{x}$ (output) from $\nu_{e}$ (input) is small. The potential possibility is that it is related to the hardness ratio of the energy spectra of $\nu_{x}$. A more detailed discussion and study on this point can be conducted after having data with richer physical details.
@@
In summary, from the heatmap computed using the KAN and SINDy-SA framework, we find that among the many input parameters determining the asymptotic values of the neutrino radial moments $I_{0}$ and $I_{1}$, the initial number density of heavy-lepton neutrinos always exhibits significant correlation. The relevance values for almost all other parameters are negligible. It is important to note that most of these parameters in fact are the initial-state combinations of the neutrino radial moments prior to the onset of FFCs. An interpretation was previously provided at the beginning of this section using the KAN model analysis, and the partial derivative method used in this section, as well as arrives at the same conclusion. This consistency may prompt us to consider the problem from a new perspective: while the triggering of FFCs is known to be dependent on the details of the initial angular distribution, the connection between the asymptotic outcome and those specific initial conditions weakens rapidly once FFCs are triggered, and the system’s dynamics become dominated by macroscopic properties such as the global particle number density. It is important to note that we do not have data on the entire evolutionary process over time, so these ideas are speculative.
@@
This behavior shares a conceptual similarity with the well-known “sandpile model” in dynamical systems exhibiting self-organized criticality [84]: the triggering of an avalanche may be caused by a single, randomly falling grain of sand, but the size, shape, and duration of the avalanche are almost entirely independent of the properties of that grain, instead being determined by the sandpile’s overall “critical state” , such as its slope and density. Once the sandpile reaches this critical state, there is no longer any correlation between the system’s response to a perturbation and the details of that perturbation itself. Similarly, in our scenario, specific initial combinations of $I_{0}$ and $I_{1}$ act as the triggering grain of sand, serving mainly to determine whether the FFC “avalanche” is initiated. Once triggered, the system’s nonlinear dynamics quickly ”forget” the details of the trigger, and the final asymptotic state is governed by some of the system’s global macroscopic properties—in this study, potentially represented by the heavy-lepton neutrino number density because the FCCs’ output energy spectrum is most sensitive to its initial density. It should be noted that this has potential consistency with the results of [65], while offering a view of the underlying dynamics.